# One-Step Offline Distillation of Diffusion-based Models via Koopman Modeling

### v2

@@ Abstract
Diffusion-based generative models have demonstrated exceptional performance, yet their iterative sampling procedures remain computationally expensive. A prominent strategy to mitigate this cost is distillation, with offline distillation offering particular advantages in terms of efficiency, modularity, and flexibility. In this work, we identify two key observations that motivate a principled distillation framework: (1) while diffusion models have been viewed through the lens of dynamical systems theory, powerful and underexplored tools can be further leveraged; and (2) diffusion models inherently impose structured, semantically coherent trajectories in latent space. Building on these observations, we introduce the Koopman Distillation Model (KDM), a novel offline distillation approach grounded in Koopman theory—a classical framework for representing nonlinear dynamics linearly in a transformed space. KDM encodes noisy inputs into an embedded space where a learned linear operator propagates them forward, followed by a decoder that reconstructs clean samples. This enables single-step generation while preserving semantic fidelity. We provide theoretical justification for our approach: (1) under mild assumptions, the learned diffusion dynamics admit a finite-dimensional Koopman representation; and (2) proximity in the Koopman latent space correlates with semantic similarity in the generated outputs, allowing for effective trajectory alignment. Empirically, KDM achieves state-of-the-art performance across standard offline distillation benchmarks—improving FID scores by up to 40% in a single generation step. All implementation details and code for the experimental setups are provided in our GitHub - https://github.com/azencot-group/KDM, or in our project page - https://sites.google.com/view/koopman-distillation-model.
@@ 1 Introduction
Diffusion-based generative models [62], including score-based and flow-matching variants [24, 41], have become ubiquitous across a wide range of domains—from images and videos to audio and time series [14, 25, 35, 12]. These models now surpass approaches such as GANs and VAEs in terms of sample quality, while also exhibiting greater training stability. Despite these advantages, one of their key limitations remains the high computational cost associated with sampling [53, 58, 45]. Generating a single high-fidelity sample typically requires executing a lengthy iterative process, progressively refining random noise through dozens or even hundreds of model evaluations [24].
@@
A widely adopted strategy to address this involves minimizing the number of inference steps [63], for example by leveraging improved numerical solvers or designing more expressive noise schedules, enabling high-quality generation with significantly fewer denoising iterations. Another increasingly popular alternative is distillation [58], where a student model learns to emulate the behavior of a teacher model. Distillation approaches vary in supervision and setup: in online distillation, the student directly learns from the teacher’s predictions during training, with the teacher providing trajectory information across time steps. In contrast, offline distillation relies on precomputed noise-image pairs generated by the teacher, and the student learns without further querying the teacher model. Offline distillation offers key advantages over online approaches: by training on precomputed noise-image pairs, it avoids the need for multi-stage training, it reduces the amount of neural function evaluations required by the teacher model, it allows to filter privacy-violating synthetic data generated by the teacher before distillation, and it eliminates the overhead of keeping the teacher model in memory throughout training [17, 19]. This leads to more private, scalable training and allows for broader reuse of distilled data across different student architectures and settings.
@@
Towards developing a principled distillation method, we identify two key observations that guide our approach: (1) While denoising-based methods leverage dynamical systems perspectives for both general generative modeling [66, 41] and distillation [78, 9], a wide range of powerful tools from dynamical systems theory remain underexplored; (2) Trained diffusion models impose a structured organization on the latent space, coherently mapping noisy inputs toward clean data. Specifically, our empirical analysis suggests that semantically related images tend to originate from nearby points in the noise space.
@@
Building on these observations, we propose leveraging a dynamical systems approach that has not yet been explored in the context of distillation: Koopman operator theory. This theory is a classical yet recently revitalized approach for representing nonlinear systems with linear dynamics in an embedded space [36, 57]. The core idea is that by mapping the system into a suitable embedding space, its evolution becomes linear. Thus, learning the diffusion process then reduces to jointly learning these embeddings and the linear operator. While Koopman operators are typically infinite-dimensional, we prove that an exact finite-dimensional representation exists under mild assumptions [26]. Moreover, prior work suggests that the Koopman operator preserves the structure inherent to certain dynamical systems [22, 6]. This motivates the use of Koopman theory to harness the observed coherent semantic structure in diffusion dynamics, offering a principled foundation for accurate teacher model distillation. To this end, we formally connect the observed structure in diffusion models to the Koopman framework by proving that this structure is preserved within the learned dynamical representation.
@@
In practice, we propose KDM (Koopman Distillation Model), a framework for modeling diffusion processes through linear evolution in a learned embedding space. Given a noisy input, an encoder maps it into an embedding space where a learned Koopman operator evolves it forward, and a decoder reconstructs the clean sample. A schematic illustration of our framework is presented in Fig. 1. The training objective encourages semantic preservation (via reconstruction loss), accurate latent linear dynamics (via Koopman consistency), and fidelity in output space (via prediction and adversarial losses). This design naturally yields fast, single-step generation, while preserving structural coherence distillation. We evaluate our method on the offline distillation benchmark, while adding two higher-resolution datasets, and achieve state-of-the-art results in both unconditional and conditional generation. Compared to prior offline work, our approach improves FID by up to 40% while requiring only a single generation step. These results demonstrate the effectiveness of our Koopman-based formulation in bridging the performance gap with online methods, offering a fast and high-fidelity offline diffusion model distillation. Our key contributions are:
@@
    Observing diffusion dynamics. We identify, analyze and formalize two key observations: (i) tools from dynamical systems theory, specifically Koopman operator theory, remain underexplored in diffusion modeling and distillation. (ii) training induces coherent semantic organization in noise space. These insights motivate a structure-promoting dynamics-aware approach to distillation.
@@
    State-of-the-art results. We conduct a comprehensive unconditional and conditional evaluation showing that our method consistently outperforms prior offline distillation approaches, achieving up to 40% FID improvement—while enabling efficient, single-step generation.
@@ 2 Related Work
Deep generative models have made significant progress in recent years, with diffusion models emerging as a leading framework for high-quality image, audio, and molecular generation. Building on principles from nonequilibrium thermodynamics and score matching [62, 66], these models define a forward process that gradually corrupts data with noise and learn to reverse it via a neural network-based denoising process. Variants like DDPMs [24] and score-based generative models [66] have shown competitive performance across a wide range of benchmarks, often surpassing VAEs and GANs [34, 21] in sample fidelity and diversity.
@@
Diffusion models generate high-quality samples but are slow due to iterative sampling. One approach improves integration using higher-order solvers and expressive noise schedules [63, 43], achieving high-quality results in significantly fewer steps. Another class of approaches seeks to distill a student model from a pre-trained teacher, reducing the sampling process to a single or few steps. Online distillation techniques [44, 58] supervise the student directly with teacher predictions during training, but require continuous access to the full teacher model, resulting in high memory and compute overhead. Recent methods demonstrate high-quality one-step generation across CIFAR-10, ImageNet, and text-to-image benchmarks [75, 81], using objectives based on distribution matching [75], score identity [81], adversarial training [59, 61], and consistency models [65, 32] among others [9, 46, 52, 82]. Offline distillation methods [19] rely on teacher-generated noise-data pairs and decouple student training from the teacher, enabling more scalable and modular pipelines. As far as we know, current distillation methods largely unexplored Koopman-based perspectives and their relation to the teacher’s underlying dynamics.
@@
Koopman operator theory provides a powerful framework for modeling nonlinear dynamical systems via linear operators acting on observable functions [36, 57]. This view has been applied to tasks such as deep learning of dynamical systems [69, 47, 73, 55], sequential disentanglement [4], and control [23]. A related strand of work focuses on identifying coherent sets—regions of the state space that evolve coherently over time—using Koopman-based embeddings [50, 18]. While these works exploit the Koopman structure for interpretability or stability, our work is, to our knowledge, the first to use Koopman-based modeling for distilling diffusion models. In this context, we demonstrate that the learned embedding space in diffusion models encodes semantically meaningful structure, and we leverage this to design a principled offline distillation framework.
@@ 3 Mathematical Background
We consider deterministic sampling in diffusion models, as introduced in the EDM framework [28]. A pre-trained diffusion model defines a time-indexed denoising vector field $f_{\theta}:\mathbb{R}^{n}\times[0,T]\rightarrow\mathbb{R}^{n}$, which maps a noisy latent state $x_{t}\in\mathbb{R}^{n}$ at time $t\in[0,T]$ toward the data manifold (at time 00). Sampling is performed by solving the reverse-time ordinary differential equation (ODE),
@@
integrated backward from a Gaussian sample $x_{T}$ at time $T$ to a clean sample $x_{0}$ at time $t=0$. The function $f_{\theta}$ is typically trained to approximate the score function $\nabla_{x_{t}}\log p_{t}(x_{t})$ or a reparameterization thereof, such as the probability flow [66]. These deterministic formulations enable sampling via high-order ODE solvers and provide a tractable setting for distillation and trajectory analysis [28].
@@
Let $\Phi_{t}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ define a (possibly time-dependent) dynamical system mapping states forward in time, i.e., $x_{t}=\Phi_{t}(x_{0})$. The Koopman operator $\mathcal{K}_{t}$ [36] is a linear operator that acts on scalar-valued observables $g:\mathbb{R}^{n}\rightarrow\mathbb{C}$ via composition,
@@
mapping each observable to its pullback along the dynamics. While the original system $\Phi_{t}$ may be nonlinear, $\mathcal{K}_{t}$ is linear on the space of observables $L^{2}(\mathbb{C}^{n})$ (with respect to a suitable measure) [16]. In classical Koopman theory, this operator is infinite-dimensional. However, if there exists a finite-dimensional subspace $\mathcal{O}=\text{span}\{g_{1},\dots,g_{d}\}\subset L^{2}(\mathbb{C}^{n})$ that is invariant under $\mathcal{K}_{t}$ for all $t\in[0,T]$, then the dynamics of $\Phi_{t}$ can be exactly represented in this lifted space via a family of linear operators $C_{t}\in\mathbb{R}^{d\times d}$, $t\in[0,T]$, such that for all $g\in\mathcal{O},\,\mathcal{K}_{t}g=C_{t}g$. This allows modeling nonlinear dynamics through linear evolution in a learned observable space, a principle that underpins our Koopman-based distillation framework.
@@ 4 Koopman Distillation of Diffusion Models
The second observation considers the structure of the trained diffusion dynamics. Score-based diffusion models learn the score $\nabla_{x_{t}}\log p_{t}(x_{t})$ of noise-corrupted data at multiple noise levels and then reverse that specific corruption process using Langevin dynamics [66] or learned reverse kernels [24] to transform Gaussian noise into images. This framework has been unified through the lens of stochastic differential equations (SDEs) by [66]. Flow matching models [41] learn a continuous velocity field that transports a simple reference distribution directly into the data distribution, where sampling is typically performed by integrating the learned ordinary differential equation (ODE). In what follows, we study the structure of the dynamical mapping arising from diffusion-based learning.
We pose the following question: Is there semantic structure in the generative process of diffusion models? In general, prior to training, there is no assurance of semantic structure between the noise and data distributions (see Fig. 2, left). However, after training, we observe the emergence of a coherent and semantically meaningful mapping from the Gaussian noise distribution to the data distribution (see Fig. 2, right). Specifically, by utilizing a simple 2D checkerboard toy dataset, we train an EDM [28] model and generate 50,000 samples, recording both the initial noise vectors and their corresponding generated outputs on the checkerboard. We then color-code each sample by its target checker cell and assign the same color to its originating noise vector. This visualization reveals that samples originating from the same semantic region in data space tend to cluster together in noise space, indicating structured and coherent dynamics. Notably, we observe a similar pattern in both flow matching model and our KDM. This emergence of coherent structure, together with prior evidence that the Koopman operator preserves structure in certain dynamical systems [22, 6], highlights the relevance of Koopman theory to diffusion dynamics. Indeed, our theoretical analysis (Sec. 5) establishes a formal link between this observed structure and Koopman theory, motivating its adoption as a principled framework for modeling these dynamics.
@@
Our Koopman distillation framework is designed by the following principle: given pairs of noisy and clean data $(x_{T},x_{0})$, where $x_{0}=\Phi_{T}(x_{T})$ for some unknown nonlinear map $\Phi_{t}$, we seek to model this transformation using a finite-dimensional linear operator in an embedding space, following Koopman theory. To achieve this, we introduce observable functions $\{g_{1},g_{2},...,g_{d}\}$, parameterized as encoder networks $E_{\phi},E_{\varphi}$ for $x_{0},x_{T}$, respectively, that map states from the original data space into a latent observable space. In this space, we assume the dynamics can be modeled linearly by a finite-dimensional Koopman operator linear layer, $C_{\eta}$. Finally, a corresponding decoder network $D_{\psi}$ maps the evolved observables back to the data space. The overall modeling objective becomes:
@@
In what follows, $d(\cdot,\cdot)$ denotes a suitable distance function, such as mean squared error (MSE) or a perceptual loss (e.g., LPIPS [77]). To effectively learn the components $E_{\phi}$, $D_{\psi}$, and $C_{\eta}$, we define a composite loss comprising three terms (see similar objectives in [47, 3]):
@@
Adversarial frameworks are widely employed in distillation methods [60, 32, 67, 68]. For example, [32] leverages an adversarial loss to enhance the fidelity of one-step generation. Motivated by these approaches, we introduce an auxiliary adversarial loss to further refine the distillation process. Specifically, we introduce a discriminator network $D_{\gamma}$, parameterized by $\gamma$, and define the adversarial loss as:
@@
where $\lambda_{\text{adv}}$ balances the adversarial component. In practice, throughout our experiments, we choose $\lambda_{\text{adv}}=0.01$. Importantly, we utilized a simple $D_{\gamma}$ with four convolutional layers. By enhancing the perceptual quality of the predictions without altering the fundamental Koopman modeling, this adversarial augmentation strengthens the overall generation performance while preserving the structured distillation at the heart of our method [5]. We present losses and more ablation studies in App. D.1. Additionally, A pseudo-code of our KDM is provided in App. E.1.
@@
A crucial aspect of generative modeling with denoising models is the ability to condition generation on auxiliary signals or labels [14]. Koopman theory provides a natural framework to incorporate such conditioning through control signals that influence the Koopman operator [8]. Formally, let $c$ denote the control signal or conditioning variable (e.g., a class label) that we want the generative process to be conditioned on. Our goal is to model the conditional dynamics $\Phi(x_{T},c)$. In the Koopman framework, we extend the observable and evolution operators to incorporate $c$, and formulate the generation as:

### v3

@@ Abstract
Diffusion-based generative models have demonstrated exceptional performance, yet their iterative sampling procedures remain computationally expensive. A prominent strategy to mitigate this cost is distillation, with offline distillation offering particular advantages in terms of efficiency, modularity, and flexibility. In this work, we identify two key observations that motivate a principled distillation framework: (1) while diffusion models have been viewed through the lens of dynamical systems theory, powerful and underexplored tools can be further leveraged; and (2) diffusion models inherently impose structured, semantically coherent trajectories in latent space. Building on these observations, we introduce the Koopman Distillation Model (KDM), a novel offline distillation approach grounded in Koopman theory—a classical framework for representing nonlinear dynamics linearly in a transformed space. KDM encodes noisy inputs into an embedded space where a learned linear operator propagates them forward, followed by a decoder that reconstructs clean samples. This enables single-step generation while preserving semantic fidelity. We provide theoretical justification for our approach: (1) under mild assumptions, the learned diffusion dynamics admit a finite-dimensional Koopman representation; and (2) proximity in the Koopman latent space correlates with semantic similarity in the generated outputs, allowing for effective trajectory alignment. KDM achieves highly competitive performance across standard offline distillation benchmarks.
@@ 1 Introduction
Diffusion-based generative models sohl2015deep , including score-based and flow-matching variants ho2020denoising ; lipman2023flow , have become ubiquitous across a wide range of domains—from images and videos to audio and time series dhariwal2021diffusion ; ho2022video ; kong2020diffwave ; coletta2023constrained ; naiman2024utilizing ; fadlon025diffusion ; gonen2025time . These models now surpass approaches such as GANs and VAEs in terms of sample quality, while also exhibiting greater training stability. Despite these advantages, one of their key limitations remains the high computational cost associated with sampling nichol2021improved ; salimans2022progressive ; luo2023latent . Generating a single high-fidelity sample typically requires executing a lengthy iterative process, progressively refining random noise through dozens or even hundreds of model evaluations ho2020denoising .
@@
A widely adopted strategy to address this involves minimizing the number of inference steps song2020denoising , for example by leveraging improved numerical solvers or designing more expressive noise schedules, enabling high-quality generation with significantly fewer denoising iterations. Another increasingly popular alternative is distillation salimans2022progressive , where a student model learns to emulate the behavior of a teacher model. Distillation approaches vary in supervision and setup: in online distillation, the student directly learns from the teacher’s predictions during training, with the teacher providing trajectory information across time steps. In contrast, offline distillation relies on precomputed noise–image pairs or teacher-generated samples only, where the student learns without any further access to the teacher model, including its weights, evaluations, or internal states.
Offline methods inherently offer key advantages over online tools. Training on precomputed, teacher-generated noise–image pairs avoids multi-stage training and repeated teacher queries, reducing both neural function evaluations and memory footprint geng2023one . It also enables pre-filtering or anonymization of potentially privacy violations fernandez2023privacy . Consequently, once the dataset is produced, the teacher can be discarded, shrinking the attack surface and supporting secure collaboration (e.g., outsourcing training without exposing proprietary information) (carlini2023extracting, ; hu2023membership, ; nasr2023scalable, ). Beyond privacy and efficiency, offline distillation is model-agnostic and architecture-general (as evidenced in our experiments), and it scales flexibly with parameter budgets. On the limitation side, offline setups require upfront generation and storage costs for the precomputed corpus and, lacking teacher access beyond the noise–image pairs, may constrain model expressivity and adaptability. By contrast, online methods may warm-start from teacher weights or query the teacher for stepwise supervision. Despite these trade-offs, offline distillation remains a practical, private, and sustainable approach to generative-model distillation. The two paradigms can be considered complementary, and the optimal choice depends on factors such as privacy requirements, computational cost, and deployment constraints.
@@
Towards developing a distillation method, we identify two key observations that guide our approach: (1) While denoising-based methods leverage dynamical systems perspectives for both general generative modeling song2021score ; lipman2023flow and distillation zhang2023fast ; chen2023geometric , a wide range of powerful tools from dynamical systems theory remain underexplored; (2) Trained diffusion models impose a structured organization on the latent space, coherently mapping noisy inputs toward clean data. Specifically, our empirical analysis suggests that semantically related images tend to originate from nearby points in the noise space.
@@
Building on these observations, we propose leveraging a dynamical systems approach that has not yet been explored in the context of distillation: Koopman operator theory. This theory is a classical yet recently revitalized approach for representing nonlinear systems with linear dynamics in an embedded space koopman1931hamiltonian ; rowley2009spectral . The core idea is that by mapping the system into a suitable embedding space, its evolution becomes linear. Thus, learning the diffusion process then reduces to jointly learning these embeddings and the linear operator. While Koopman operators are typically infinite-dimensional, we prove that an exact finite-dimensional representation exists under mild assumptions iacob2023finite . Moreover, prior work suggests that the Koopman operator preserves the structure inherent to certain dynamical systems goyal2025deep ; boulle2024multiplicative . This motivates the use of Koopman theory to harness the observed coherent semantic structure in diffusion dynamics, offering a principled foundation for accurate teacher model distillation. To this end, we formally connect the observed structure in diffusion models to the Koopman framework by proving that this structure is preserved within the learned dynamical representation.
@@
In practice, we propose KDM (Koopman Distillation Model), a framework for modeling diffusion processes through linear evolution in a learned embedding space. Given a noisy input, an encoder maps it into an embedding space where a learned Koopman operator evolves it forward, and a decoder reconstructs the clean sample. A schematic illustration of our framework is presented in Fig. 1. The training objective encourages semantic preservation (via reconstruction loss), accurate latent linear dynamics (via Koopman consistency), and fidelity in output space (via prediction and adversarial losses). This design naturally yields fast, single-step generation, while preserving structural coherence distillation. We evaluate our method on the offline distillation benchmark, while adding two higher-resolution datasets, and achieve highly competitive results in both unconditional and conditional generation. These results demonstrate the effectiveness of our Koopman-based formulation in bridging the performance gap with online methods, offering a fast and high-fidelity offline diffusion model distillation. Our key contributions are:
@@
    Observing diffusion dynamics. We identify, analyze and formalize two key observations: (i) tools from dynamical systems theory, specifically Koopman operator theory, remain underexplored in diffusion modeling and distillation; (ii) training induces coherent semantic organization in noise space. These insights motivate a structure-promoting dynamics-aware approach to distillation.
@@
    Highly competitive results. We conduct a comprehensive unconditional and conditional evaluation showing that our method outperforms prior offline distillation approaches, achieving FID improvement, while enabling efficient, single-step generation.
@@ 2 Related Work
Deep generative models have made significant progress in recent years, with diffusion models emerging as a leading framework for high-quality image, audio, and molecular generation. Building on principles from nonequilibrium thermodynamics and score matching sohl2015deep ; song2021score , these models define a forward process that gradually corrupts data with noise and learn to reverse it via a neural network-based denoising process. Variants like DDPMs ho2020denoising and score-based generative models song2021score have shown competitive performance across a wide range of benchmarks, often surpassing VAEs and GANs kingma2014auto ; goodfellow2014generative in sample fidelity and diversity.
@@
Diffusion models generate high-quality samples but are slow due to iterative sampling. One approach improves integration using higher-order solvers and expressive noise schedules song2020denoising ; lu2022dpm , achieving high-quality results in significantly fewer steps. Another class of approaches seeks to distill a student model from a pre-trained teacher, reducing the sampling process to a single or few steps. Online distillation techniques luhman2021knowledge ; salimans2022progressive supervise the student directly with teacher predictions during training, but require continuous access to the full teacher model, resulting in high memory and compute overhead. Recent methods demonstrate high-quality one-step generation across CIFAR-10, ImageNet, and text-to-image benchmarks yin2024one ; zhou2024score , using objectives based on distribution matching yin2024one , score identity zhou2024score , adversarial training sauer2024fast ; sauer2024adversarial , and consistency models song2023consistency ; kim2024consistency among others chen2023geometric ; luo2023diff ; nguyen2024swiftbrush ; zhou2024simple ; berman2025said . Offline distillation methods geng2023one rely on teacher-generated noise-data pairs and decouple student training from the teacher. Aside from the concurrent work turan2025unfolding , current distillation methods largely unexplored Koopman-based perspectives and their relation to the teacher’s underlying dynamics.
@@
Koopman operator theory provides a powerful framework for modeling nonlinear dynamical systems via linear operators acting on observable functions koopman1931hamiltonian ; rowley2009spectral . This view has been applied to tasks such as deep learning of dynamical systems takeishi2017learning ; lusch2018deep ; yeung2019learning ; otto2019linearly ; naiman2023operator ; naiman2024generative , sequential disentanglement berman2023multifactor ; barami2025disentanglement , and control han2022desko . A related strand of work focuses on identifying coherent sets—regions of the state space that evolve coherently over time—using Koopman-based embeddings mezic2005spectral ; froyland2010coherent . While these works exploit the Koopman structure for interpretability or stability, our work focuses on using Koopman-based modeling for distilling diffusion models. In this context, we demonstrate that the learned embedding space in diffusion models encodes semantically meaningful structure, and we leverage this to design a principled offline distillation framework.
@@ 3 Mathematical Background
We consider deterministic sampling in diffusion models, as introduced in the EDM framework karras2022elucidating . A pre-trained diffusion model defines a time-indexed denoising vector field $f_{\theta}:\mathbb{R}^{n}\times[0,T]\rightarrow\mathbb{R}^{n}$, which maps a noisy latent state $x_{t}\in\mathbb{R}^{n}$ at time $t\in[0,T]$ toward the data manifold (at time 0). Sampling is performed by solving the reverse-time ordinary differential equation (ODE),
@@
integrated backward from a Gaussian sample $x_{T}$ at time $T$ to a clean sample $x_{0}$ at time $t=0$. The function $f_{\theta}$ is typically trained to approximate the score function $\nabla_{x_{t}}\log p_{t}(x_{t})$ or a reparameterization thereof, such as the probability flow song2021score . These deterministic formulations enable sampling via high-order ODE solvers and provide a tractable setting for distillation and trajectory analysis karras2022elucidating .
@@
Let $\Phi_{t}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ define a (possibly time-dependent) dynamical system mapping states forward in time, i.e., $x_{t}=\Phi_{t}(x_{0})$. The Koopman operator $\mathcal{K}_{t}$ koopman1931hamiltonian is a linear operator that acts on scalar-valued observables $g:\mathbb{R}^{n}\rightarrow\mathbb{C}$ via composition,
@@
mapping each observable to its pullback along the dynamics. While the original system $\Phi_{t}$ may be nonlinear, $\mathcal{K}_{t}$ is linear on the space of observables $L^{2}(\mathbb{C}^{n})$ (with respect to a suitable measure) eisner2015operator . In classical Koopman theory, this operator is infinite-dimensional. However, if there exists a finite-dimensional subspace $\mathcal{O}=\text{span}\{g_{1},\dots,g_{d}\}\subset L^{2}(\mathbb{C}^{n})$ that is invariant under $\mathcal{K}_{t}$ for all $t\in[0,T]$, then the dynamics of $\Phi_{t}$ can be exactly represented in this lifted space via a family of linear operators $C_{t}\in\mathbb{R}^{d\times d}$, $t\in[0,T]$, such that for all $g\in\mathcal{O},\,\mathcal{K}_{t}g=C_{t}g$. This allows modeling nonlinear dynamics through linear evolution in a learned observable space, a principle that underpins our Koopman-based distillation framework.
@@ 4 Koopman Distillation of Diffusion Models
The second observation considers the structure of the trained diffusion dynamics. Score‑based diffusion models learn the score $\nabla_{x_{t}}\log p_{t}(x_{t})$ of noise‑corrupted data at multiple noise levels and then reverse that specific corruption process using Langevin dynamics song2021score or learned reverse kernels ho2020denoising to transform Gaussian noise into images. This framework has been unified through the lens of stochastic differential equations (SDEs) by song2021score . Flow matching models lipman2023flow learn a continuous velocity field that transports a simple reference distribution directly into the data distribution, where sampling is typically performed by integrating the learned ordinary differential equation (ODE). In what follows, we study the structure of the dynamical mapping arising from diffusion-based learning.
We pose the following question: Is there semantic structure in the generative process of diffusion models? In general, prior to training, there is no assurance of semantic structure between the noise and data distributions (see Fig. 2, left). However, after training, we observe the emergence of a coherent and semantically meaningful mapping from the Gaussian noise distribution to the data distribution (see Fig. 2, right). Specifically, by utilizing a simple 2D checkerboard toy dataset, we train an EDM karras2022elucidating model and generate 50,000 samples, recording both the initial noise vectors and their corresponding generated outputs on the checkerboard. We then color-code each sample by its target checker cell and assign the same color to its originating noise vector. This visualization reveals that samples originating from the same semantic region in data space tend to cluster together in noise space, indicating structured and coherent dynamics. Notably, we observe a similar pattern in both flow matching model and our KDM. This emergence of coherent structure, together with prior evidence that the Koopman operator preserves structure in certain dynamical systems goyal2025deep ; boulle2024multiplicative , highlights the relevance of Koopman theory to diffusion dynamics. Indeed, our theoretical analysis (Sec. 5) establishes a formal link between this observed structure and Koopman theory, motivating its adoption as a principled framework for modeling these dynamics.
@@
Our Koopman distillation framework is designed by the following principle: given pairs of noisy and clean data $(x_{T},x_{0})$, where $x_{0}=\Phi_{T}^{-1}(x_{T})$ for some unknown nonlinear map $\Phi_{t}^{-1}$, we seek to model this transformation using a finite-dimensional linear operator in an embedding space, following Koopman theory. To achieve this, we introduce observable functions $\{g_{1},g_{2},...,g_{d}\}$, parameterized as encoder networks $E_{\phi},E_{\varphi}$ for $x_{0},x_{T}$, respectively, that map states from the original data space into a latent observable space. In this space, we assume the dynamics can be modeled linearly by a finite-dimensional Koopman operator linear layer, $C_{\eta}$. Finally, a corresponding decoder network $D_{\psi}$ maps the evolved observables back to the data space. The overall modeling objective becomes:
@@
In what follows, $d(\cdot,\cdot)$ denotes a suitable distance function, such as mean squared error (MSE) or a perceptual loss (e.g., LPIPS zhang2018unreasonable ). To effectively learn the components $E_{\phi}$, $D_{\psi}$, and $C_{\eta}$, we define a composite loss comprising three terms (see similar objectives in lusch2018deep ; azencot2020forecasting ):
@@
Adversarial frameworks are widely employed in distillation methods sauer2023adversarial ; kim2024consistency ; ladd2024latent ; sdxllightning2024progressive . For example, kim2024consistency leverages an adversarial loss to enhance the fidelity of one-step generation. Motivated by these approaches, we introduce an auxiliary adversarial loss to further refine the distillation process. Specifically, we introduce a discriminator network $D_{\gamma}$, parameterized by $\gamma$, and define the adversarial loss as:
@@
where $\lambda_{\text{adv}}$ balances the adversarial component. In practice, throughout our experiments, we choose $\lambda_{\text{adv}}=0.01$. Importantly, we utilized a simple $D_{\gamma}$ with four convolutional layers. By enhancing the perceptual quality of the predictions without altering the fundamental Koopman modeling, this adversarial augmentation strengthens the overall generation performance while preserving the structured distillation at the heart of our method blau2018perception . We present losses and more ablation studies in App. D.1. Additionally, A pseudo-code of our KDM is provided in App. E.1.
@@
A crucial aspect of generative modeling with denoising models is the ability to condition generation on auxiliary signals or labels dhariwal2021diffusion . Koopman theory provides a natural framework to incorporate such conditioning through control signals that influence the Koopman operator azencot2013operator ; brunton2022modern . Formally, let $c$ denote the control signal or conditioning variable (e.g., a class label) that we want the generative process to be conditioned on. Our goal is to model the conditional dynamics $\Phi(x_{T},c)$. In the Koopman framework, we extend the observable and evolution operators to incorporate $c$, and formulate the generation as: